Three balls of masses m1,m2 and m3 are lying in a straight line. the first ball is moved with a certain velocity so that it strikes the second ball directly and itself comes to rest. The second ball collides with the third and is itself reduced to rest. If e be the coefficient of restitution for each ball then m1/m2 is same as??


a) m2/m3 


b)m2/2m3


c)m2.m3


d)none

Grade:11

1 Answers

Ashwin Muralidharan IIT Madras
12 years ago

Hi Kritika,

 

Use conservation of momentum.

 

m1v1 = m2v2

and m2v2 = m3v3

 

Also by writing the eqn for co-eff of restitution:

v2 = ev1

and v3 = ev2.

 

So m1/m2 = v2/v1 = e = v3/v2 = m2/m3.

 

So option (A).

 

Hope that helps.

 

Best Regards,

Ashwin (IIT Madras).